OpenStudy (anonymous):

Solve: e^3x = 8 A. ln 5 B. 3 ln 8 C. ln 8/3 D. ln 3/8

OpenStudy (anonymous):

@saifoo.khan Can you help?

OpenStudy (saifoo.khan):

\[3x = \ln 8\]

OpenStudy (saifoo.khan):

Can you solve now?

OpenStudy (anonymous):

Then I divide by 3x?? Making the answer be C??

OpenStudy (saifoo.khan):

\[\frac{\ln8}{3}\]

OpenStudy (anonymous):

Thanks! But what about: 1/e^x = 7

OpenStudy (anonymous):

\[\frac{1}{e^x}=7\iff e^x=\frac{1}{7}\]in equivalent exponential form this means \[x=\ln(\frac{1}{7})\]

OpenStudy (anonymous):

or you can write \[\frac{1}{e^x}=e^{-x}\] so that \[-x=\ln(7)\] or \[x=-\ln(7)\]

OpenStudy (anonymous):

answers are the same becauese \[\ln(\frac{1}{7})=\ln(7^{-1})=-\ln(7)\]

OpenStudy (anonymous):

Thanks :)